If the sum of two numbers is 50 and their mean proportional is 15, then the numbers are-
Answer by mananth(16946)

If the sum of two numbers is 50 and their mean proportional is 15, then the numbers are-Let one number be x
The other number then will be 50-x
mean proportion is15
(50-x)/15 = 15/x
x(50-x) = 225
50x-x^2=225
x^2-50x+225=0
(x-45)(x-5)=0
5 and 45 are the numbers